OpenStudy (anonymous):

A right triangle has a hypotenuse of 25 feet and a leg of 7 feet. Which of these is the length of the other leg? A. 12 Feet B. 15 Feet C. 21 Feet D. 24 Feet

OpenStudy (anonymous):

Pythagoras theorem says the hypotenuse squared is equal to the squares of the other two sides added. 25^2=7^2 + x^2 x^2=576 x=24 Answer is D.

OpenStudy (anonymous):

Thankss ! :)